OpenStudy (anonymous):

Identify intervals on which the function is increasing, decreasing, or constant. g(x) = 2 - (x - 7)2 Increasing: x < 2; decreasing: x > 2 Increasing: x < -7; decreasing: x > -7 Increasing: x < 7; decreasing: x > 7 Increasing: x > 2; decreasing: x < 2

OpenStudy (jdoe0001):

http://nayak2010.files.wordpress.com/2011/01/picture-of-vertex-form-of-parabola.png <--- vertex form of a parabola

OpenStudy (jdoe0001):

keep in mind that \(\bf g(x) = 2 - (x - 7)^2\implies g(x) = - (x - 7)^2+2\) at the vertex is where the parabola changes from decreasing to increasing or the viceversa

OpenStudy (anonymous):

ok so how do I start to solve this problem?

OpenStudy (jdoe0001):

well... what do you think would be the vertex for it?

OpenStudy (anonymous):

i know that to find vertex of something you can graph it and look at the peak of the parbala and that's the vertex

OpenStudy (anonymous):

and this one is just a straight line so its endless

OpenStudy (jdoe0001):

right... well is not a straight line, \(\bf g(x) = 2 - (x - 7)^2\)

OpenStudy (jdoe0001):

unless you meant \(\bf g(x) = 2 - (x - 7)2\)

OpenStudy (anonymous):

yes sir that's what i grapghed your last oe you typed

OpenStudy (anonymous):

ok so the vertex is 2 correct

OpenStudy (jdoe0001):



OpenStudy (anonymous):

oh so its how do i know which > < it is? i know its going to be 2 though just not sure which > < it is

OpenStudy (jdoe0001):

a function is increasing when it greater as it moves towards the right-side and decreases as it's lesser as it moves towards the right-side

OpenStudy (anonymous):

Increasing: x > 2; decreasing: x < 2 Increasing: x < 2; decreasing: x > 2 how do i tell which one of these it is they both say increasing and decreasing they just have different greater than or less than signs

OpenStudy (jdoe0001):

2-(x-7)2 => 16-x

OpenStudy (anonymous):

-_- im sorry what

OpenStudy (jdoe0001):

so, let us graph that... .let's say use 2 points y = 16-x x = 0 y = 16-0 => y = 16 so ( 0, 16 ) x =10 y = 16-10 => y = 6 so ( 10, 6) 

OpenStudy (anonymous):

so its this one Increasing: x < 2; decreasing: x > 2

OpenStudy (jdoe0001):

well... notice, if it's a straight line, it's just decreasing

OpenStudy (jdoe0001):

but I'd think it's a parabola

OpenStudy (anonymous):

like those are my answers and all of them both have increasing and decreasing on them

OpenStudy (jdoe0001):

yeap, meaning is \(\bf g(x) = 2 - (x - 7)^2\)

OpenStudy (anonymous):

so im right?